### v1

# DRPruning: Efficient Large Language Model Pruning through Distributionally Robust Optimization

Abstract
Large language models (LLMs) deliver impressive results but face challenges from increasing model sizes and computational costs. Structured pruning reduces model size and speeds up inference but often causes uneven degradation across domains, leading to biased performance. To address this, we propose DRPruning, which incorporates distributionally robust optimization to restore balanced performance across domains, along with further improvements to enhance robustness. Experiments in monolingual and multilingual settings show that our method surpasses similarly sized models in pruning and continued pretraining over perplexity, downstream tasks, and instruction tuning. We further provide analysis demonstrating the robustness of our method towards various domains and distribution shifts. Furthermore, our method automatically determines optimal reference losses and data ratios, suggesting potential for broader applications. Our code is available at https://github.com/hexuandeng/DRPruning.
DRPruning: Efficient Large Language Model Pruning through Distributionally Robust Optimization
Hexuan Deng1    Wenxiang Jiao    Xuebo Liu1thanks:   Corresponding Author    Min Zhang1    Zhaopeng Tu 1 Institute of Computing and Intelligence, Harbin Institute of Technology, Shenzhen, China {hxuandeng,wenxiangjiaonju,tuzhaopeng}@gmail.com, {liuxuebo,zhangmin2021}@hit.edu.cn
1 Introduction
Large language models (LLMs) have advanced rapidly, achieving impressive results across a wide range of tasks (Bang et al., 2023; Jiao et al., 2023; Frieder et al., 2023; Bian et al., 2024). However, this progress has come with increasing model sizes, significantly raising computational costs for both training and inference, which impacts their accessibility. Structured pruning is a promising approach to reduce model size (Han et al., 2015; Wen et al., 2016), but it often causes uneven performance degradation across domains, leading to biased capabilities and unfair downstream task performance (Xia et al., 2024).
To address this, we propose a distributionally robust pruning method that dynamically adjusts the data distribution during training to ensure balanced performance across domains. Building on distributionally robust optimization (DRO; Oren et al., 2019; Sagawa et al., 2019), which aims to optimize worst-case performance across distributions, a reference loss is defined for each domain as a target. A higher weight is then assigned to domains with larger deviations from this reference loss, while not straying too far from a predefined reference data ratio. However, setting these hyperparameters is challenging, and suboptimal configurations often result in poor outcomes (Zhou et al., 2021).
To resolve this, we introduce a method that dynamically adjusts hyperparameters during training. Domains with worse performance often exhibit higher losses, deviating further from the expected loss after sufficient training. Using scaling laws (Kaplan et al., 2020; Ghorbani et al., 2022), we predict the loss after training as the reference, where larger deviations indicate poorer performance, thereby promoting capability recovery in these areas. Additionally, we gradually increase the reference data ratio for domains with greater deviations, ensuring robustness across a wider range of distributions, particularly more challenging ones.
We validate our approach through experiments in monolingual and multilingual settings, which represent varying degrees of distributional shift. Our model consistently outperforms Xia et al. (2024) and other LLMs in both pruning and continued pretraining, as measured by perplexity (-5.59% and -5.60%, respectively), downstream tasks (+1.27% and +1.78%, respectively), and instruction tuning (55.4% win rate). Particularly in multilingual settings, our model achieves +2.95% in downstream tasks. To further assess domain-specific performance, we develop a sentence continuation benchmark using existing unlabeled data. Results show strong consistency with downstream task performance, confirming the benchmark’s reliability and further demonstrating our improved domain-level capabilities (+17.9% on average). Moreover, our approach to determining reference losses and data ratios could potentially enhance data efficiency across a broader range of applications, driving further advancements in LLMs.
Our contributions are summarized as follows:
  • We tackle domain imbalance in structured pruning by introducing a distributionally robust pruning method, that dynamically adjusts data ratios during training to ensure robustness against distributional shifts.
  • We demonstrate the method’s effectiveness through extensive experiments in monolingual and multilingual settings. Our method surpasses existing ones in perplexity, downstream tasks, instruction tuning, and domain-specific evaluation.
  • We offer refined reference losses and data ratios for training, which can be applied more broadly to enhance various model training processes and contribute to advancements for LLMs.
2 Background
2.1 Structured Pruning
To prune the model to any target configuration, we adopt structured pruning based on Sheared Llama (Xia et al., 2024). For each granularity $i$, we learn pruning masks $Z=\{\mathbf{z}^{i}\mid\mathbf{z}^{i}\in\mathbb{R}^{D_{i}}\}$ that determine whether substructures are pruned or retained, where $z^{i}_{j}=0$ indicates pruning of the $j$-th substructure. We apply pruning across granularities that include transformer layers, hidden dimensions, attention heads, and FFN intermediate dimensions.
To parameterize the masks, we use the $\ell_{0}$ regularization method (Louizos et al., 2018) with hard concrete distributions that concentrate probability mass at 0 or 1. We further use Lagrange multipliers to ensure the pruned model meets the target configuration. Specifically, if exactly $t^{i}$ parameters must be retained for $\mathbf{z}^{i}$, we impose the constraint:
~$\tilde{\ell}^{i}=\lambda^{i}\left(\sum_{j}z^{i}_{j}-t^{i}\right)+\phi^{i}\left(\sum_{j}z^{i}_{j}-t^{i}\right)^{2}.$ (1)
The final training loss integrates these constraints with the language modeling loss of the pruned model, jointly optimizing the model parameters $\theta$ and pruning masks $\mathbf{z}$, where $\mathbf{z}$ typically uses a higher learning rate. After pruning, we retain the highest-scoring components.
2.2 Distributionally Robust Optimization
To mitigate uneven domain performance after pruning, we apply distributionally robust optimization (DRO; Oren et al., 2019; Sagawa et al., 2019) to improve the model’s robustness to distribution shifts. DRO seeks a model $\theta$ that performs well across a set of potential test distributions $\mathcal{Q}$ over $n$ domains. Formally, we aim to:
[…]
To solve the min-max optimization, we use the iterative best response (IBR) algorithm (Fudenberg and Levine, 1998). Each iteration consists of first performing Empirical Risk Minimization (ERM) on the current data distribution $\mathbf{q}^{t}$, followed by updating the data distribution using worst-case weights based on the current parameters. Formally,
[…]
$\mathbf{q}^{t+1}\leftarrow\underset{\mathbf{q}=\{q_{1},\ldots,q_{n}\}\in\mathcal{Q}}{\operatorname{argmax}}\sum_{i}q_{i}\ell\left(\theta^{t+1};D_{i}\right).$
3 Methodology
In Section 3.1, we introduce our overall procedure to utilize these components to design an effective pruning method, as shown in Figure 1. Further, to determine the data ratio $\mathbf{q}^{t}$, DRO requires two inputs: loss function $\ell\left(\theta;D\right)$ and potential distributions $\mathcal{Q}$, which strongly influence $\mathbf{q}^{t}$ and, consequently, the model’s training process. Therefore, we discuss how to better calculate these parameters in Section 3.2 and 3.3.
Figure 1: Data proportion update procedure for DRPruning. The gray part represents the standard training process, the yellow part represents the normal process for DRO, and the blue part represents our newly added module.
3.1 Distributionally Robust Pruning
Building on the aforementioned methods, we construct a pruning approach to enhance the model’s robustness to both capacity reduction and distribution shifts. Specifically, we adopt common techniques (Ma et al., 2023; Li et al., 2024; Zhang et al., 2024b), first applying structured pruning to reduce model parameters, followed by continued pretraining to restore capabilities. Compared to training from scratch, this approach requires only a small amount of unlabeled data to restore the model’s performance (Xia et al., 2024).
Combining DRO and pruning.
We further integrate the DRO method into both the pruning and continued pretraining stages. During training, DRO dynamically adjusts the data ratio to improve the model’s robustness and convergence speed. Specifically, to prevent overfitting, we compile a validation set and use the evaluation loss as the loss score. After each evaluation (every 50 steps during pruning and every 400 steps during continued pretraining), we update the data ratio based on the evaluation loss using the DRO method, directing training to focus more on underperforming domains.
Further improvement.
Next, we optimize the loss function (Section 3.2) and potential distributions (Section 3.3) to ensure robust training, as shown in Figure 1. In contrast, Sheared Llama employs a dynamic scheduling strategy that forces the model to strictly adhere to the relative loss magnitudes of larger LLMs, without placing any constraints on the potential distributions. This leads to suboptimal results, particularly in multilingual settings with significant distribution shifts.
3.2 Dynamic Loss Function
To stabilize DRO training and prevent domains with slow convergence from disproportionately influencing the weights, the use of a reference loss $\ell_{R}$ is a common approach (Oren et al., 2019; Zhou et al., 2021). This reference loss establishes the minimum acceptable performance for a domain. Furthermore, we update the loss score as $\ell\left(\theta;D\right)\leftarrow\ell\left(\theta;D\right)-\ell_{R}$. Proper tuning of $\ell_{R}$ can significantly improve performance (Jiao et al., 2022). However, determining an appropriate value remains a challenging task.
Minimum performance estimation.
To address this, we predict the model’s loss at the end of training as an estimate of the minimum acceptable performance. Specifically, we leverage scaling laws to capture training dynamics and forecast the loss based on evaluation loss trends (Kaplan et al., 2020; Zhang et al., 2024a). Given the number of parameters $P$ and the current training step $T$, the predicted training loss is estimated by:
^$\hat{\ell}(P,T)=A\cdot\frac{1}{P^{\alpha}}\cdot\frac{1}{T^{\beta}}+E,$ (4)
where $A$, $E$, $\alpha$, and $\beta$ are trainable parameters. For each domain, after each evaluation, we collect a data point, refit the curve to all collected points, and use the predicted curve to estimate the loss at the end of training as the predicted minimum performance. Following Hoffmann et al. (2022a), we estimate using the Huber loss ($\delta$ = 0.001) and the L-BFGS algorithm, and select the average of the best-fitting three from a grid of initializations. To ensure sufficient data points, we start predictions only after 20% of training is complete.
Reference loss adjustment.
Subsequently, we set the reference loss using the predicted minimum performance. In our preliminary experiments, this approach exhibits strong numerical stability. To accelerate convergence, we adopt the minimum value as the reference loss. This dynamically evaluates domains with poorer performance, allowing DRO to assign higher weights to these domains, thereby promoting faster model convergence.
3.3 Dynamic Potential Distribution
Sagawa et al. (2019) consider robustness to arbitrary subpopulations, which is overly conservative and degenerates into training only on the highest-loss domain. To address this issue, Zhou et al. (2021) propose a more reasonable assumption by restricting $\mathcal{Q}$ in Equation 2 to an $f$-divergence ball (Csiszár, 1967) around a reference data ratio $\mathbf{p}_{R}$. This yields promising results, better ensuring domain balance (Jiao et al., 2022). Formally,
$\mathcal{Q}=\left\{\mathbf{q}:\chi^{2}\left(\mathbf{q},\mathbf{p}_{R}\right)\leq\rho\right\}.$ (5)
However, this assumption can be too restrictive, necessitating a carefully chosen reference data ratio $\mathbf{p}_{R}$. An unreasonable choice may reduce the model’s robustness to distributional shifts.
Reference data ratio adjustment.
To address this, we propose a method that combines the strengths of the aforementioned approaches. We still employ Equation 5 to constrain the distribution within a limited range, while gradually shifting the reference data ratio towards domains with higher losses to improve the model’s robustness to more challenging distributions. To ensure adequate training across all traversed potential distributions, we gradually update the reference ratio.
Compared to existing reference ratios, the DRO method dynamically assigns higher weights $\mathbf{q}$ to domains with higher losses. This method shows good numerical stability, which we leverage to update the reference ratio. Formally, we update:
$\mathbf{p}^{t+1}_{R}=\delta\cdot\mathbf{q}^{t}+(1-\delta)\cdot\mathbf{p}^{t}_{R}.$ (6)
Finally, to prevent the method from degenerating into training solely on the highest-loss domain, we constrain the reference ratio of each domain to lie between $\frac{1}{n}$ and $n$ times the initial ratio. Formally, we set $\frac{1}{n}\cdot\mathbf{p}^{0}_{R}\leq\mathbf{p}^{t}_{R}\leq n\cdot\mathbf{p}^{0}_{R}$. We apply this method after 40% of the training is completed, ensuring the model sufficiently converges near the initial reference ratio and that the reference loss stabilizes.
Overall process.
As illustrated in Figure 1, we treat each evaluation phase as an iteration. First, the evaluation loss is used to update the reference loss. Then, using the new reference loss along with the previous reference data ratio as input to the DRO process, we obtain a new data proportion for the next training step and update the reference data ratio accordingly.
4 Experiments
4.1 Experimental Setup
To ensure comparability with Sheared Llama, we maintain consistency with most of its experimental settings. Further details of our experimental setup are provided in Appendix A.1.
Data.
We allocate 0.4 billion tokens for pruning, utilizing the publicly available pruning dataset of Sheared Llama. We employ 50 billion tokens for continued pretraining. To enhance model quality, we use SlimPajama (Shen et al., 2023), a filtered version of RedPajama (Computer, 2023), and use its training split for continued pretraining.
Model configurations.
We use the Llama2-7B model (Touvron et al., 2023b) as the base model throughout our main experiments. We employ the same target architecture as Sheared Llama for structured pruning to ensure a fair comparison. We compare our results to strong open-source models of similar sizes, including Pythia-1.4B and 2.8B (Biderman et al., 2023) and Sheared Llama-1.3B and 2.7B (Xia et al., 2024). Additionally, we reproduce Sheared Llama, using the same data settings to control for other variables (ReSheared).
Downstream task evaluation.
We use the lm-evaluation-harness package (Gao et al., 2024) to evaluate on an extensive suite of downstream tasks:
  • We follow Llama2 to report the 0-shot performance on PIQA (Bisk et al., 2020), WinoGrande (WinoG, Sakaguchi et al., 2020), ARC Easy (ARCE, Clark et al., 2018), SQuAD (Rajpurkar et al., 2018), BoolQ (Clark et al., 2019), TruthfulQA (TruthQA, Lin et al., 2022), and 5-shot performance on Natural Questions (NQ, Kwiatkowski et al., 2019) and TriviaQA (TriQA, Joshi et al., 2017).
  • We follow Pythia to report the 0-shot performance of LAMBADA (LAMB, Paperno et al., 2016), LogiQA (Liu et al., 2020), SciQ (Welbl et al., 2017), and WSC (Kocijan et al., 2020).
  • We follow Sheared Llama to report performance of the tasks used by Open LLM Leaderboard, including 10-shot HellaSwag (HelS, Zellers et al., 2019), 25-shot ARC Challenge (ARCC, Clark et al., 2018), and 5-shot MMLU (Hendrycks et al., 2021).
Instruction tuning evaluation.
To further explore the potential applications of the base model, we follow Sheared Llama by training with 10,000 instruction-response pairs sampled from the ShareGPT dataset and using another 1,000 instructions for evaluation. We follow Wang et al. (2024) and Xia et al. (2024) to employ LLMs, specifically GPT-4o, as an evaluator to compare the responses of the two models and report the win rates.
Method From To PPL $\downarrow$ Down. $\uparrow$
Sheared Llama 7B 1.3B 10.05 34.89
ReSheared 7B 1.3B 10.42 34.85
DRPruning 7B 1.3B 9.83 35.60
Sheared Llama 7B 2.7B 7.64 39.75
ReSheared 7B 2.7B 7.83 39.98
DRPruning 7B 2.7B 7.40 40.18
Table 1: Perplexity (PPL) and downstream task performance (Down.) of pruned models. “Sheared Llama” is the open-source model, “ReSheared” is our reproduction of Sheared Llama, and “DRPruning” refers to the method proposed in this paper. “From” and “To” denote model size before and after pruning.
Figure 2: The curve of PPL changes during pruning from 7B for our reproduction of Sheared Llama and our method. Over the first 640 iterations (the vertical dash line), the model size is gradually reduced from 7B to the target size, which causes an initial increase in PPL.
Tasks 7B 2.7B 1.3B
Llama2† Pythia† Sheared† ReSheared DRPrun. Pythia† Sheared† ReSheared DRPrun.
WSC 36.54 38.46 48.08 36.54 46.15 36.54 36.54 40.38 50.00
TriQA (5) 64.16 27.17 42.92 40.14 43.33 18.19 26.03 24.98 28.10
NQ (5) 25.98 7.12 14.85 13.49 15.82 4.79 8.75 8.39 10.44
TruthQA 32.09 28.79 30.21 28.41 30.13 30.75 29.12 28.09 29.68
LogiQA 30.11 28.11 28.26 26.27 28.73 27.50 27.50 28.11 28.88
BoolQ 77.71 64.50 65.99 64.92 65.08 63.30 62.05 61.01 63.36
LAMB 73.90 64.76 68.21 66.18 66.91 61.67 61.09 58.84 60.28
MMLU (5) 44.18 27.09 26.63 25.70 26.99 26.75 25.70 26.60 27.28
SciQ 94.00 88.50 91.10 90.10 89.80 86.70 87.00 86.40 87.70
ARCE 76.35 64.27 67.34 67.72 67.13 60.40 60.90 60.35 60.90
ARCC (25) 52.65 36.35 42.66 40.10 40.53 33.02 33.96 34.30 33.62
PIQA 78.07 73.88 76.12 76.71 75.19 70.84 73.50 74.59 72.69
WinoG 69.06 59.83 65.04 63.38 64.72 57.38 57.85 60.06 58.01
SQuAD 40.02 26.81 49.26 49.17 44.69 22.66 29.57 37.59 35.06
HelS (10) 78.95 60.81 71.24 72.03 69.22 53.49 61.05 63.06 58.88
Average 58.25 46.43 52.53 50.72 51.63 43.60 45.37 46.18 46.99
Table 2: Our method (DRPrun.) outperforms publicly available models of comparable size on downstream tasks, as well as the Sheared Llama strategies. “Sheared” is the open-source model, while “ReSheared” is our reproduction. The number of shots is indicated in parentheses, with 0-shot used if unspecified. A model marked with †  denotes pretraining on different data. ReSheared and our method share the same experimental setup.
4.2 Main Results for Pruning
Our method surpasses the Sheared Llama during pruning in PPL and downstream tasks.
Our method significantly accelerates convergence.
We record the average perplexity (PPL) across different domains on the validation set. Table 1 shows that after 3.2k steps, PPL drops by 2.60% compared to the open-source model and by 5.59% in a fair comparison. Figure 2 confirms a more stable PPL as pruning proceeds, with further potential in later training stages, suggesting additional gains with extended training.
Our method better preserves the original model’s performance during pruning.  For a comprehensive analysis, we evaluate the model’s downstream task performance post-pruning. Across 15 tasks, our method consistently improves results, with an average gain of +1.53% over the open-source model and +1.27% in a fair comparison.
The pruning similarity analysis is detailed in Appendix B.1, and Appendix B.2 shows that pruning larger LLMs offers no advantage.
4.3 Main Results for Continued Pretraining
We show that our method outperforms existing LLMs and other pruning approaches on both standard LM benchmarks and instruction tuning.
The LLMs recovered by our method are better foundation models after continued pretraining.
Table 2 presents the downstream performance of models with similar sizes, showing that our model surpasses most open-source models. Although we cannot replicate the 2.7B Sheared Llama results due to data differences, our model still outperforms most open-source LLMs, achieving +4.95% on average. Additionally, under consistent experimental conditions, our method outperforms our reproduction of Sheared Llama, achieving +1.78% on average. Moreover, we achieve significantly lower perplexity: 5.61 vs 5.97 for 1.3B, and 5.00 vs 5.27 for 2.7B, averaging a -5.60% reduction.
The effectiveness of our method is further demonstrated by supervised finetuning.
We compare the win rates of our instruction-tuned model with Sheared Llama. Figure 3 shows our model achieves a 55.4% win rate, surpassing both the open-source Sheared Llama and our reproduction. This highlights our model offers a stronger foundation for further use.
Figure 3: Win rate during instruction tuning. “Sheared Llama” refers to the open-source version, while “ReSheared” denotes our reproduced version. Our approach consistently outperforms Sheared Llama across all settings.
Figure 4: Effectiveness of our method compared to constant scheduling and naïve DRO during the 1.3B continued pretraining. The left figure shows PPL trends and the right shows average performance across 15 downstream tasks.
Figure 5: Variation of data ratio, reference data ratio, and reference loss during continued pretraining in the main experiment. The top three plots show results for pruning to 1.3B parameters, while the bottom three are for pruning to 2.7B parameters. Dashed lines in the data ratio plots represent the initial reference data ratio for each domain. Reference loss plots display the difference from the initial value, with the dashed line at $y=0$ indicating no change from the initial reference loss.
5 Analysis
5.1 Ablation Study
We assess the impact of our dynamic data scheduling on 1.3B models, comparing it to constant scheduling and the naïve DRO method.
Our method significantly outperforms the DRO method.
As shown in Figure 4, our method consistently reduces PPL compared to the constant approach (-0.067 by the end of training), while the DRO method shows a slight increase (+0.011). In downstream tasks, our approach also outperforms both baselines, with performance improving steadily in the mid to late stages of training (-0.26 for DRO, +0.95 for ours). This underscores the sensitivity of DRO to hyperparameters and the necessity of dynamic adjustments.
CC C4 GitHub Book Wiki ArXiv StackEx
Data Ratio Constant 67.0% 15.0% 4.5% 4.5% 4.5% 2.5% 2.0%
DRO 54.6% 29.0% 3.7% 5.0% 3.8% 1.9% 1.9%
DRPruning 55.7% 15.6% 2.4% 3.7% 18.4% 2.1% 2.1%
Reference Ratio vs. Constant 47.1% 30.5% 2.9% 3.5% 9.1% 3.9% 3.0%
Loss Evaluation Loss vs. Constant +0.011 -0.010 +0.030 +0.007 -0.123 +0.003 +0.001
Reference Loss vs. Constant +0.067 +0.162 +0.093 +0.094 -0.108 +0.014 -0.053
Table 3: Data usage ratios, hyperparameter adjustments, and domain-specific loss for Constant, naïve DRO, and our strategy. The first three rows show the total data usage ratios during training, while the last three rows represent the hyperparameters and the evaluation loss at the end of training.
Our method dynamically identifies underperforming domains.
To highlight how our method differs from DRO, Table 3 presents data usage ratios, hyperparameter adjustments, and domain-specific loss. Our strategy identifies Wiki as underperforming compared to optimal, assigning it a lower reference loss (-0.11) and a higher reference ratio (+4.6%). This reduces loss on Wiki (-0.12) while keeping loss stable in other domains (maximum increase of +0.03). Besides, by dynamically adjusting the reference ratio, we improve the model’s robustness across the full range of ratios from the initial to the new reference distribution. Compared to DRO, this offers robustness to a wider range of distribution shifts.
Furthermore, we analyze how our strategy dynamically adjusts training parameters, as illustrated in Figure 5. Both the 1.3B and 2.7B models show similar trends, increasing allocation to C4 and Wiki data and reduced allocation to CC, suggesting that our method provides reliable training expectations. This observation also underscores the limitations of static hyperparameter settings. The high potential of Wiki improvement leads to a significantly increased reference ratio, while in later training stages, slower loss convergence in the CC domain prompts a reduced weight.
Our method enables effective dynamic evaluation of reference loss.
To accelerate convergence, we select the minimum predicted value as the reference loss, which raises concerns about the inability to increase the reference loss when domain potential decreases. However, the two rightmost figures show that our predictions are conservative and consistently decrease during training. When potential declines, the rate of decrease slows, resulting in a relatively higher reference loss. This outcome stems from using the loss of a limited training step rather than the fully trained loss in scaling laws, resulting in cautious estimates.
Figure 6: The generation procedure for our sentence continuation task. The orange nodes represent data storage nodes, while the blue trapezoidal nodes represent data processing nodes.
Model CC C4 GitHub Book Wiki ArXiv StackExchange Average
1.3B Constant 35.00 40.00 88.00 47.75 21.75 82.00 72.50 55.29
Sheared Llama 21.75 26.00 93.75 33.50 29.50 38.50 56.50 42.79
ReSheared 30.50 30.25 89.25 32.00 23.00 81.00 47.50 47.64
DRPruning 44.00 51.50 94.75 48.00 33.50 86.50 90.00 64.04
2.7B Sheared Llama 81.25 89.50 95.50 96.50 89.25 90.50 82.75 89.32
ReSheared 61.75 60.25 96.00 73.00 80.50 93.75 92.00 79.61
DRPruning 82.25 77.75 99.00 86.75 87.50 79.50 89.25 86.00
Table 4: Domain-level results under the benchmark we generated. The abbreviations CC, …, Stackexchange refer to the seven domains used for training in RedPajama.
5.2 Robustness across Different Domains
While our method shows clear advantages in PPL, a gap remains between PPL and downstream performance. The lack of domain-specific test sets also limits further analysis. To address this, we create downstream tasks from unlabeled datasets for detailed domain-specific evaluation.
Automatic construction of sentence continuation tasks across domains.
To assess base model performance, we use sentence continuation tasks where the model selects the better continuation between two options. As shown in Figure 6, we use the SlimPajama test set as the correct sentence and have the model generate the incorrect alternatives. First, “LLM Filtering” selects consecutive sentences with a causal relationship, then “Split” each sentence into two parts: the first as the question, and the second as the right option. For this large-scale filtering, we use GPT-4o-mini.
Next, “LLM Rewrite” generates incorrect options. Since LLMs struggle to create incorrect but related content, we follow Deng et al. (2024) to first generate a reasonable continuation, then modify it to create an incorrect version. Finally, “LLMs Verification” scores both options and filters out cases where the scores don’t match the true answers. We use GPT-4o in these procedures. We further select questions with the largest score differences. In total, we select 400 questions per domain. A case is shown in Table 5.
Question If the latter described their efforts to adapt to European conditions,
Right Option the former insisted that Muslims adhere to proper canons of learning and textual interpretation.
Wrong Option it also highlighted the resilience and ingenuity that had brought them this far despite challenges.
Table 5: Case For the sentence continuation task.
Our method outperforms the baselines consistently across the domains.
As shown in Table 4, our method consistently outperforms the Constant and Sheared Llama scheduling strategies, achieving better downstream performance in most domains. This demonstrates that our approach not only improves learning in hard domains but also preserves or enhances performance in others. Additionally, our benchmark results align well with the average performance across 15 tasks, performing slightly below the open-source Sheared Llama 2.7B model. This consistency confirms the quality and reliability of our benchmark and results.
Base Model Prune PT Method EN RU ZH JA AR TR KO TH Average
XGLM-1.7B X X - 55.06 52.97 51.02 51.00 42.89 37.99 49.00 38.63 47.32
Qwen1.5-1.8B X X - 60.89 52.30 56.13 53.30 42.17 34.98 48.25 36.75 48.10
Qwen2-1.5B X X - 61.58 57.83 55.72 55.30 43.31 35.98 49.25 36.02 49.37
Qwen2-1.5B X $\checkmark$ ReSheared 62.16 58.95 54.93 55.60 43.91 37.27 54.05 39.96 50.85
Qwen2-1.5B X $\checkmark$ DRPruning 61.67 59.09 54.01 54.95 45.14 46.91 52.65 44.42 52.35
Qwen2-7B $\checkmark$ $\checkmark$ DRPruning 60.43 56.80 55.72 55.05 45.69 43.82 53.95 43.53 51.87
Table 6: The average performance on downstream tasks across multiple languages. “Prune” refers to the pruning procedure applied to the base model, while “PT” indicates continued pretraining on the provided dataset.
EN RU ZH JA AR TR KO TH
Data Ratio Default Reference Ratio 27.7% 18.5% 13.0% 10.4% 9.1% 8.9% 6.7% 5.8%
ReSheared 82.8% 5.9% 5.5% 2.2% 1.0% 0.8% 1.2% 0.6%
DRPruning 19.0% 7.8% 12.9% 19.1% 9.2% 19.9% 6.7% 5.4%
Reference Ratio vs. ReSheared 23.1% 9.2% 8.5% 20.3% 8.6% 17.8% 6.7% 5.8%
Loss Evaluation Loss vs. ReSheared +0.143 -0.067 -0.123 -0.330 -0.465 -0.841 -0.282 -0.304
Reference Loss vs. ReSheared +0.211 +0.008 -0.060 -0.267 -0.412 -0.804 -0.219 -0.229
Table 7: Data usage ratios, hyperparameter adjustments, and domain-specific loss for reproduction of Sheared Llama (ReSheared) and our approach (DRPruning) during continued pretraining from Qwen2 1.5B.
5.3 Robustness under Distribution Shifts
To verify our method under larger distribution shifts, we conduct experiments under the multilingual setting.
Data setups.
We utilize the CulturaX dataset (Nguyen et al., 2024), a comprehensive multilingual resource that integrates mC4 and OSCAR, and has been rigorously cleaned and deduplicated. For our analysis, we focus on eight languages supported by Qwen2: English (EN), Russian (RU), Chinese (ZH), Japanese (JA), Arabic (AR), Turkish (TR), Korean (KO), and Thai (TH), which represent diverse language families.
Experimental setups.
We use Qwen2 series models (Yang et al., 2024) as the base models and explore two approaches: (1) continued pretraining from Qwen2-1.5B, and (2) pruning Qwen2-7B and then continued pretraining. Due to grouped query attention differences, we keep the head dimension unchanged, resulting in a 1.8B target architecture. The reference loss is initialized with Qwen2-7B on the validation set. For the reference ratio, we follow Conneau et al. (2020), upsampling low-resource languages with a smoothing rate of 0.3. We select various downstream tasks and report average performance. Detailed model configurations and metrics are provided in Appendix A.2.
Our method demonstrates superior distribution robustness compared to Sheared Llama.
As shown in Table 6, our method outperforms Sheared Llama with an average gain of +1.50, including +3.48 in low-resource languages. It also surpasses the open-source Qwen2-1.5B, with an average gain of +2.98, and +6.14 in low-resource languages.
Our method dynamically allocates higher weights to low-resource languages.
As shown in Table 7, Sheared Llama focuses on hard-to-improve areas like English, where its loss remains above the reference loss, leading to a highly imbalanced data distribution. In contrast, our method dynamically identifies underperforming domains, increasing the reference loss for high-resource languages and lowering it for low-resource ones. This leads to more balanced data scheduling and better evaluation loss, further demonstrating the robustness of our approach in handling distribution shifts.
Continued pretraining from the pruned model underperforms from the pre-trained ones.
Continued pretraining from the pruned model result in a slight performance drop (average -0.48), despite the increased parameters (1.8B vs. 1.5B). This contrasts with Xia et al. (2024), where continued pretraining on a smaller model shows minimal gains. In our case, using different data and a stronger small model improves performance during continued pretraining, leading to different results.
6 Related Work
Pruning.
Unstructured pruning (Frankle and Carbin, 2019; Frantar and Alistarh, 2023; Sun et al., 2024) removes individual weights but offers limited speedup. This study focuses on structured pruning (Han et al., 2015; Wen et al., 2016), which removes entire structural components, such as neurons or layers, making it more effective for simplifying models and improving efficiency. In task-specific models, extensive pruning can retain performance (Liu et al., 2017; Wang et al., 2020a; Lagunas et al., 2021; Xia et al., 2022; Kurtic et al., 2023). However, for large language models (LLMs), as the ratio of training data to model size increases (Hoffmann et al., 2022b; Touvron et al., 2023a, b), fewer redundant parameters remain, leading to significant performance degradation after pruning.
To counter this, performance recovery techniques like continued pretraining are essential (Ma et al., 2023; Zhang et al., 2024b). However, continued pretraining of pruned models reduces loss at different rates across domains compared to training from scratch, resulting in less efficient data utilization (Xia et al., 2024). To address this, we employ a distributionally robust pruning method that dynamically adjusts the data distribution during training, ensuring balanced performance across domains.
Distributionally robust optimization (DRO).
Overparameterized neural networks often perform well on i.i.d. test sets but struggle with underrepresented data groups (Hovy and Søgaard, 2015; Blodgett et al., 2016; Tatman, 2017; Duchi et al., 2020). To address this, unlike empirical risk minimization (ERM), which minimizes expected loss for a fixed distribution, MultiDDS (Wang et al., 2020b) uses a gradient-based meta-learning method to optimize the sampling distribution. However, MultiDDS increases computational and memory costs. In contrast, DRO (Delage and Ye, 2010; Ben-Tal et al., 2012; Bertsimas et al., 2014) improves performance without additional complexity (Hashimoto et al., 2018; Oren et al., 2019; Levy et al., 2020).
DRO finds a model that performs well across multiple possible test distributions. Group DRO (Sagawa et al., 2019) minimizes the worst-case loss over all domains without constraining potential distribution, while CVaR-Group DRO (Oren et al., 2019) averages the largest $N$ group losses. These methods can be overly conservative, as they account for robustness to arbitrary subpopulations. Zhou et al. (2021) address this by constraining potential distribution within an $f$-divergence ball (Csiszár, 1967) around a reference data ratio, yielding promising results (Jiao et al., 2022).
DRO enhancement.
DRO shows strong performance but relies on two main hyperparameters. The first is the reference loss, usually set by training an additional baseline model (Zhou et al., 2021; Jiao et al., 2022), though this is expensive for LLMs. Xia et al. (2024) attempt to reduce costs by using scaling laws to predict the pruned model’s performance, but our experiments show this method fails under severe distribution shifts during continued pretraining. The second hyperparameter is the reference data ratio, often determined through temperature-based sampling (Arivazhagan et al., 2019; Conneau et al., 2020) or manually (Touvron et al., 2023a; Parmar et al., 2024). However, fixed ratios can hinder model convergence in challenging distributions. We propose a dynamic method that shifts weight toward higher-loss domains, enhancing distribution robustness and improving downstream performance.
7 Conclusion
We present DRPruning, a distributionally robust pruning method that addresses uneven performance degradation across domains during structured pruning. By utilizing and further improving distributionally robust optimization (DRO), our method focuses more on domains with poorer performance, accelerating performance recovery. It outperforms existing models and training methods in both monolingual and multilingual settings, achieving lower perplexity, higher task accuracy, and better instruction tuning. Further analysis demonstrates the robustness of our method against various domains and distribution shifts. Additionally, the dynamic adjustment of reference loss and data ratios exhibits broad applicability, with strong potential to support balanced training across diverse tasks.
Limitations
Exploration of smaller pruning ratios.
Due to computational constraints, we are unable to explore pruning to larger models, i.e., employing smaller pruning ratios. Retaining a larger proportion of the model’s parameters may lead to different outcomes in some experiments. For example, it remains to be investigated whether pruning larger models provides benefits, and whether it is better to continue pretraining from a pruned model or from a smaller, fully pre-trained model.
More extensive continued pretraining.
Xia et al. (2024) point out that pruned models exhibit higher training ceilings. Although good performance can be achieved with tens of billions of training samples, this study does not investigate whether training the models to full convergence using hundreds or thousands of billions of samples would yield better results than continuing pretraining from existing pre-trained models under similar settings.
Validation in other scenarios.
We have validated our method’s effectiveness in the pruning phase, the pruning recovery phase, and the continued pretraining phase. However, our method is expected to be applicable in broader contexts, such as pretraining from scratch and cross-domain instruction tuning. Broader validation would further demonstrate the superiority of our approach.
Ethics Statement
Our work adheres to the ACL Ethics Policy and uses publicly available datasets for reproducibility. LLMs may exhibit racial and gender biases, so we strongly recommend users assess potential biases before applying the models in specific contexts. Additionally, due to the difficulty of controlling LLM outputs, users should be cautious of issues arising from hallucinations.
Appendix A Detailed Experimental Setup
Pruning Contined Pretraining
Training Steps 3,200 48,000
Learning rate of $z,\phi,\lambda$ $1.0$
Learning Rate of $\theta$ $0.0001$ $0.0001$
LR warmup ratio $10\%$ $3\%$
Batch size (tokens) $131$K $1$M
Ratio update interval $m$ (steps) $50$ $400$
Table 8: Training hyperparameters for the main experiment.
Model #Param #Layers Hidden Intermediate #Heads Head Dim
Pruned-1.3B 1.3B 24 2048 5504 16 128
Pruned-2.7B 2.7B 32 2560 6912 20 128
Llama2-7B 6.7B 32 4096 11008 32 128
Table 9: The model configurations for the target model of pruning and the base models for the main experiment.
A.1 Main Experiment
Model training.
All experiments are conducted on 8 NVIDIA A100 40GB GPUs. The training hyperparameters for the main experiment are listed in Table 8, and the target model configuration for pruning is detailed in Table 9. Pruning takes approximately 12 hours for both the 1.3B and 2.7B models. Continued pretraining requires around 9 days for the 1.3B model and 18 days for the 2.7B model.
For both pruning and continued pretraining, we follow the configurations of Sheared Llama as closely as possible. We use fully sharded data parallel (Zhao et al., 2023) for parallel training and FlashAttention V1 (Dao et al., 2022) to speed up the training process. A cosine learning rate scheduler is employed, reducing the learning rate to 10% of its peak value.
DRO.
For the DRO setup, we follow Zhou et al. (2021) closely. The constraint size $\rho$ for the chi-square ball is set to {0.05, 0.1, 0.2}. Preliminary experiments show that $\rho=0.1$ yields the best results, so we use this value in all experiments. Following their setup, we truncate the dynamic data ratio to prevent it from dropping below the minimum reference data ratio, which further ensures balanced domain training. We compute historical loss values using an exponential moving average, with the hyperparameter $\lambda$ set to 0.1, which is also used for updating the reference data ratio. Besides, for the prediction of the reference loss, we maintain an average loss below $3\times 10^{-5}$, demonstrating the effectiveness of our method.
Instruction tuning.
For instruction tuning, the instruction begins with "You are a helpful assistant. Write a response that appropriately completes the request." We perform full-parameter fine-tuning for 5 epochs, with a learning rate of $5e^{-5}$, a warmup ratio of $3\%$, and a batch size of $128$.
To evaluate instruction tuning, we follow the methodology of Sheared Llama, using LLMs to assess model performance. Given outputs from two models, we ask the LLM to determine which is better using the prompt: “Here is the user request: …. Here are the two outputs for this request: Output A: …. Output B: …. Which output is better, A or B?”. Since Wang et al. (2024) note that using GPT models as evaluators can lead to preference shifts when output order is reversed, we randomly switch the positions of the outputs to ensure each result appears as Output A or Output B equally. We report the average win rate to mitigate position bias. The model gpt-4o-2024-08-06 is used for evaluation.
Model #Param #Layers Hidden Intermediate #Heads #KV Heads Head Dim
Pruned-1.8B 1.8B 28 1536 8960 14 2 128
Qwen2-1.5B 1.5B 28 1536 8960 12 2 128
Qwen2-7B 7.6B 28 3584 18944 28 4 128
Table 10: The model configurations for the target model of pruning and the base models for the multilingual experiment.
A.2 Multilingual Experiment
Model training.
The experimental setup is largely consistent with Appendix A.1. Given the extended training duration, we standardize the training to 40,000 steps. The configuration of the target model for pruning is detailed in Table 10. To maintain a consistent ratio between the number of heads and KV heads required by the structured pruning method, we keep the head dimension unchanged and add two extra heads, increasing the number of parameters to 1.8B. Continued pretraining takes around 8 days for the 1.5B model and 12 days for the 1.8B model.
Metrics.
We adopt the experimental setups from previous studies and evaluate performance on downstream tasks in a zero-shot setting. Specifically, we follow XGLM (Lin et al., 2021) and mGPT (Shliazhko et al., 2024), covering tasks such as natural language inference (XNLI; Conneau et al., 2018), Winograd schema challenge (XWINO; Tikhonov and Ryabinin, 2021), commonsense reasoning (XStoryCloze; Lin et al., 2021), and paraphrase detection (PAWSX; Yang et al., 2019). Task coverage varies across languages, and not all tasks include all languages in our training set. We report results for languages overlapping between tasks and our training set, providing average performance if a language appears in multiple tasks. The lm-evaluation-harness package (Gao et al., 2024) is used for the comprehensive evaluation of downstream tasks.
Appendix B Supplement Experimental Results
B.1 Analysis of Mask Similarity
Figure 7: Convergence of masks over 3200 pruning steps. Similarity indicates the similarity between pruning decisions at a certain step and the final decisions at step 3200. “Layer”, “Hidden”, “Head”, and “Intermediate” correspond to the four pruning dimensions.
Structure Same Different
Layer 81.25$\pm$3.61 78.65$\pm$2.24
Hidden 60.33$\pm$0.78 65.02$\pm$1.47
Head 68.88$\pm$0.47 70.53$\pm$0.65
Intermediate 56.37$\pm$0.15 56.40$\pm$0.14
Table 11: Mask similarity mean values and the Standard Error of the Mean (SEM) under different data scheduling strategies and random seeds. “Same” indicates using identical data scheduling but different random seeds, while “Different” indicates using different data scheduling strategies.
Tasks Pruning Continued Pretraining
To: 1.3B To: 2.7B To: 1.3B To: 2.7B
From: 7B From: 13B From: 7B From: 13B From: 7B From: 13B From: 7B From: 13B
ARCC (25) 23.21 22.10 30.29 27.30 33.62 32.17 40.53 40.36
ARCE 42.26 40.07 53.11 47.31 60.90 58.92 67.13 66.58
BoolQ 59.69 59.88 59.36 60.12 63.36 56.88 65.08 67.13
HelS (10) 35.27 32.38 48.07 41.62 58.88 58.70 69.22 67.67
LAMB 38.27 34.08 51.80 46.56 60.28 59.87 66.91 66.23
LogiQA 26.73 25.50 27.19 24.42 28.88 25.65 28.73 29.80
MMLU (5) 24.82 25.78 24.86 25.56 27.28 26.76 26.99 27.60
NQ (5) 1.91 1.52 4.02 2.35 10.44 8.86 15.82 13.24
PIQA 61.81 61.32 67.08 64.15 72.69 72.31 75.19 74.27
SciQ 79.80 79.60 86.30 83.30 87.70 87.30 89.80 91.00
SQuAD 13.80 6.87 17.05 18.61 35.06 28.52 44.69 46.55
TriQA (5) 5.26 3.33 11.75 6.77 28.10 24.40 43.33 38.17
TruthQA 32.99 34.15 31.12 31.50 29.68 29.66 30.13 30.75
WinoG 51.62 49.09 54.14 53.83 58.01 56.83 64.72 62.04
WSC 36.54 36.54 36.54 36.54 50.00 60.58 46.15 36.54
Average 35.60 34.15 40.18 38.00 46.99 45.83 51.63 50.53
Table 12: The performance of pruning Llama2-7B and 13B models down to 1.3B and 2.7B parameters. “Pruning” refers to using the pruned model without continued pretraining, while “Continue Pretraining” means using the model after continued pretraining. “From” and “To” indicate the size of the source and target model, respectively.
To perform a detailed analysis of the masks, we extract the masks generated during training and prune the model by removing components with the lowest scores, shaping the model according to the target specifications. We then calculate the probability of the model making consistent pruning decisions for each substructure and examine how different training steps or strategies influence the masks. Specifically, we apply Sheared Llama, constant, and our proposed strategies, using two distinct random seeds for pruning, which yields six unique 1.3B models. We then analyze the similarities across these models.
The masks converge quickly during training.
The convergence speed of the masks during training is illustrated in Figure 7. Pruning achieves over 75% similarity within 400 steps and over 95% within 800 steps, indicating that effective results can be obtained with relatively few pruning steps. While layer pruning converges rapidly, pruning intermediates of fully connected layers is slower, suggesting that coarser-grained decisions converge more quickly than finer-grained decisions.
The randomness of pruning decisions is significant.
We analyze mask similarity across training sessions with different random seeds under identical and distinct data scheduling strategies. The results, shown in Table 11, present mean values and Standard Error of the Mean (SEM). The trend across pruning dimensions aligns with the previous findings: similarity is higher at the coarse-grained layer level and lower at the finer-grained intermediate level. Besides, comparisons across three pairs under identical settings and twelve pairs under different settings show consistently low similarity. However, no significant difference in perplexity is observed, suggesting that the model’s interchangeable parameters allow similar outcomes despite different pruning decisions. This randomness obscures variations caused by differing data distributions.
B.2 Pruning from Larger LLMs
We investigate whether pruning from larger language models (LLMs) with a higher pruning ratio provides additional benefits. Experiments are conducted in the monolingual setting, consistent with the main text, to compare the effects of pruning from Llama2-7B and Llama2-13B.
The results, presented in Table 12, indicate that pruning from the 13B model consistently yields worse outcomes, regardless of whether continued pretraining is applied. On average, this approach results in a downstream performance decrease of 1.47. These findings suggest that pruning from a larger model leads to a more significant performance decline, often producing inferior results under a fixed training budget, especially under a high pruning ratio.